A zero-coupon bond with face value $1,000 and maturity of 3 years sells for $950.0. What is its yield to maturity? Enter your answer as a decimal, rounded to four decimal places Your Answer: Answer
Answer #1

In zero coupon bond is a bond but does not pay interest on bond by company or others. Instead of paying interest, it sell to investor less than the face value or equal to face value on maturity the company will pay face value or above the face value respectively.

YTm for zero coupon bond=(F/PV)1/n - 1

F= face value or maturity value

PV= present value or purchase price

n= number of periods

Here F = 1000 PV= 950. n =3   

YTM= (1,000÷950)1/3 - 1

=0.017447. ( in decimals)

= 1.7447%. ( in Percentage)